b.The time it takes to multiply two matrices of size n is determined by the polynomial expression 2n³ – n². For example, it takes 45 operations to multiply two 3x3 matrices. Compute the number of operations required to run a program that multiplies two matrices of size 20.

Answer by Edwin McCravy(20060)

Since n = 20, just substitute in

2n³ – n²

But since you asked, you must not know how to do that, so:

Everywhere you see an n in 2n³ – n², write (20) instead, like this: 

2(20)³ - 20²

Next do the exponents (20)³, which means to multiply 

20×20×20 = 400×20 = 8000

Also, we do the other exponent (20)² which means to multiply 20×20 = 400

So in

2(20)³ - 20²

we write (8000) in place of the (20)³ and 400 in place of the 20²

2(8000) - 400

Then we multiply 2×8000 and get 16000.  So we write 16000 in place of
the 2(8000)

16000 - 400

Finally we subtract and get

15600

So it takes 15,600 operations to multiply two matrices of size 20. 

Edwin